For what values of x and y does (x+y)^3=x^3+y^3?
Answer by stanbon(75887)

(x+y)^3 = x^3+y^3
x^3+3x^2y+3xy^2+y^3=x^3+y^3
3x^2y+3xy^2=0
3xy(y+x)=0
x=0,or y=0, or y=-x
Cheers,
Stan H.